OpenStudy (anonymous):

[4.06] Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4). y = -2x + 5 y = 2x - 3 y = 2x + 6 y = -2x + 2

OpenStudy (anonymous):

@Heather_94 @hartnn

OpenStudy (anonymous):

@sauravshakya @satellite73 @sasogeek

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Help please I think it is the last one

OpenStudy (anonymous):

@mattfeury

OpenStudy (anonymous):

@TheWriter

OpenStudy (anonymous):

@estudier

OpenStudy (sasogeek):

do you know how to find the slope when 2 points are given?

OpenStudy (anonymous):

Yeah the slope is 2

OpenStudy (anonymous):

I already did that

OpenStudy (anonymous):

is the answer the last one

OpenStudy (anonymous):

@hashsam1

OpenStudy (anonymous):

HELPPPPPPPPPPPPPPPPP

mathslover (mathslover):

You can verify it @owenmisa123 , did you try to do that?

mathslover (mathslover):

put the values in the fourth option that you got (y=-2x+1) of : i) y = 6 and x = -2 y = -2x + 2 ii) y = 4 and x = -1 y =-2x + 2

mathslover (mathslover):

if you get : LHS + RHS , then you are right *(LHS --> left hand side and RHS --> Right Hand Side)

OpenStudy (sasogeek):

\( \frac{\huge y-4}{\huge x+1}=\huge 2 \) solve for y

OpenStudy (anonymous):

thank you guise

OpenStudy (anonymous):

The slope is -2, watch your signs

OpenStudy (anonymous):

formula of slope = y2-y1/x2-x1 4-6/-1-(-2) = -2/+1 = -2 y=mx+x = general equation of line select any one point first and find value of C 4=-1*2 +c 4=-2+c 6=c now put value of c and m y=mx+c y=+2x+6